article spadre2008 graphicx 000 000 Title of the Article (Sometimes Extending over Two or Author One^{1} and Author Two^{2,a} This is a template article, to be used along with the style spadre2008.sty, prepared for manuscripts to

the Proceedings of the 24th Winter Workshop on Nuclear Dynamics, South Padre, Texas, April 5–12, 2008. Articles so prepared are suitable for submissions to e.g. the LANL preprint server as well.

list of keywords, relevant to the article specifications see, e.g. http://www.aip.org/pacs/ document page1 Introductionintro A historical, methodological or logical scope outlined in the Introduction helps the readers place the sub-

ject of the article or it enables to guess the present stage of the subject. Ample citation bib1,bib2,bib3,bib4,bib5,bib6,bib7,bib8,bib9 of literature that deals with the subject in question is also given in this section, as a source of information. In a second part, the Introduction tells the readers more details about the structure of the article. It often serves as a guide to introduce the sections that build up the article.

At the beginning of the template file spadre2008-template.tex, there is a header in which certain data are asked to submit for communication purposes.

Technical–Preparatory Sectiontechno The technical or methodological preparatory section of the article should commence the row of the

sections. The experimental description of measurements or the mathematical introduction of the problem, the definitions of the concepts or the descriptions of the methods, procedures applied in the investigations should be treated in this section to start with the problem bib1,bib6,bib7,bib8,bib9.

Citation of references mentioned in the article will be done as seen in the previous sentence. Citations referring to some journals should be given, for example, as in Ref. bib4, using the ‘protocol’ shown in the References.

Subsections dealing with detailsdetails Subsections (maybe, the deeper \subsubsections) will be arranged so that they appear as discernible

distinct parts within the whole structure of the article. For too wide tables or figures it may happen that a “landscape page” will be setup.

Mathematicsmaths Mathematics can be e.g. embedded P_{1}(p) = E_{p}(dN/d^{3}p) as in this case, or displayed separately from

text rows as (the boldface letter p denotes a vector): equation n_{μ}(x) = ∫_{Σ} d^{3}σ_{μ}(x′) δ^{(4)}(x − x′) .eq1